Consider the following information for Maynor Company, which uses a periodic inventory system Transaction Beginning InventoryRequired A Required BRequired C FIFO Ending Inventory Cost of Goods SoldRequired A Required B E Required C LIFO Ending Inventory Cost of Goods SoldRequired A Required B Required C Weighted Average (Round the per unit cost to two decimal places and then round your answers
Answer #1

FIFO method states that inventory purchased first is sold first.

Under periodic inventory method, the records are updated at the end of the period and not continuously

Units in ending inventory = 87-29-24

= 34 units

Ending inventory = 29*78 + 5*72

= $2,622

Cost of goods sold = Total Cost – Cost of Ending Inventory

= $6,230-2,622

= $3,608

b.LIFO method states that inventory purchased later will be sold first

Ending Inventory = 12*62 + 22*68

= $2,240

Cost of Goods Sold = 6,230-2,240 = $3,990

C,Weighted average method takes into account weighted average cost

Weighted average cost = 6,230/87 = $71.61

Ending inventory = 34*71.61 = $2,435

Cost of goods sold = 3,795